Investment Return MedTech Corp stock was $50.95 per share at the end of last year. Since then, it paid a $0.45 per share dividend. The stock price is currently $62.50. If you owned 500 shares of MedTech, what was your percent return?

Multiple Choice

  • 7.20 percent

  • 19.20 percent

  • 23.55 percent

  • 8.83 percent
Answer #1
P0 = Share price at the end of last year = 50.95
P1 = Share price at present = 62.50
D = Dividend = 0.45
Percent return = ( P1 - P0 + D ) / P0 = ( 62.50 - 50.95 + 0.45 ) / 50.95 23.55%
Answer : 23.55 percent